OpenStudy (anonymous):

The sum of two numbers is 48 and the difference is 6. what are the numbers? @niguyver

OpenStudy (anonymous):

x+y=48 x-y=6 Solving by elimination: x+y=48 x-y=6 ====== 2x = 54 x = 27 Substituting x back in: 27-y = 6, therefore y = 21 27 and 21

OpenStudy (campbell_st):

a + b = 48 a - b = 6 solve simultaneously just add the equations to eliminate b

OpenStudy (hockeychick23):

x+y=48 x-6=y subtract the first equation from the second. y+6= 48-y add y to both sides 2y+6= 48 subtract 6 from both sides 2y= 42 y= 21 x= y+6 x= 27

OpenStudy (anonymous):

thanks guys I found the answer